OpenStudy (anonymous):

how do you Write the equation of each circle. Center(3,7) with a radius of 4 units.

OpenStudy (anonymous):

(x-3)^2+(y-7)^2=16

OpenStudy (anonymous):

what does the radius do or do i not use it?

OpenStudy (anonymous):

radius is 4^2=16

OpenStudy (anonymous):

general equation of a circle is \[(x-a)^{2}+(y-b)^{2} = r^{2}\] a and b are the center. and r is the radius.